OpenStudy (anonymous):

Can you solve and show your steps so I can see how to solve these types of questions? There are 52 cards in a deck. There are 4 eights and 4 nines. What is the probability that if you cut the deck, you'll get an eight or a nine? Express your answer as a decimal number rounded to the hundredths place.

jimthompson5910 (jim_thompson5910):

how many 8s or 9s are there?

OpenStudy (anonymous):

4..

jimthompson5910 (jim_thompson5910):

there are 4+4 = 8 cards that are either 8 or 9

jimthompson5910 (jim_thompson5910):

this is out of 52 cards total

OpenStudy (anonymous):

reqd. p\[reqd. probability=\frac{ 4 }{ 52 }+\frac{ 4 }{52 }\]

jimthompson5910 (jim_thompson5910):

so the probability as a fraction is 8/52 = 2/13 use a calculator to get 2/13 = 0.15384615 and this rounds to 0.15, which is your final answer

OpenStudy (anonymous):

Ohhhhhh ok, thanks sooooo much :)

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

hey jim could you help me

jimthompson5910 (jim_thompson5910):

yeah either send me a direct message or tag me

OpenStudy (anonymous):

\[P \left( A \cup B \right)=P \left( A \right)+P \left( B \right)\] \[P \left( A \right)=\frac{ 4 }{ 52 }\] Similarly P(B)